In right triangle ABC, b=7 and c=14. Find the missing side, a.
Answer by swazzie.(1)

USE PYTHAGOREAN THEOREM : a^2 + b^2 = c^2
a^2 + 7^2 = 14^2
a^2 + 49 = 196
a^2 = 196 - 49
a^2 = 149
a^2= 12.21